# : A Unified Large Vision-Language Model for Remote Sensing Visual Grounding

### v2

@@ Abstract
Remote sensing (RS) visual grounding aims to use natural language expression to locate specific objects (in the form of the bounding box or segmentation mask) in RS images, enhancing human interaction with intelligent RS interpretation systems. Early research in this area was primarily based on horizontal bounding boxes (HBBs), but as more diverse RS datasets have become available, tasks involving oriented bounding boxes (OBBs) and segmentation masks have emerged. In practical applications, different targets require different grounding types: HBB can localize an object’s position, OBB provides its orientation, and mask depicts its shape. However, existing specialized methods are typically tailored to a single type of RS visual grounding task and are hard to generalize across tasks. In contrast, large vision-language models (VLMs) exhibit powerful multi-task learning capabilities but struggle to handle dense prediction tasks like segmentation. This paper proposes GeoGround, a novel framework that unifies support for HBB, OBB, and mask RS visual grounding tasks, allowing flexible output selection. Rather than customizing the architecture of VLM, our work aims to elegantly support pixel-level visual grounding output through the Text-Mask technique. We define prompt-assisted and geometry-guided learning to enhance consistency across different signals. To support model training, we present refGeo, a large-scale RS visual instruction-following dataset containing 161k image-text pairs. Experimental results show that GeoGround demonstrates strong performance across four RS visual grounding tasks, matching or surpassing the performance of specialized methods on multiple benchmarks. Code available at https://github.com/zytx121/GeoGround.
@@ 1 Introduction
To support GeoGround training and promote the development of visual RS grounding, we introduce refGeo, a large-scale RS visual grounding instruction-following dataset. refGeo consolidates four existing visual grounding datasets from RS [26, 37, 9, 13] and introduces a new aerial vehicle visual grounding dataset (AVVG). AVVG extends traditional 2D visual grounding to a 3D context, enabling VLMs to perceive 3D space from 2D aerial imagery. For each referred object, we provide HBB, OBB, and mask, with the latter automatically generated by the SAM [8].
@@
We introduce refGeo, the largest RS visual grounding instruction-following dataset, consisting of 161k image-text pairs and 80k RS images, including a new 3D-aware aerial vehicle visual grounding dataset.
@@ 2 Related Work
Compared to multimodal tasks like image captioning [40, 14], text-image retrieval [19], and visual question answering (VQA) [17] in RS, research on referring detection is a novel task with limited research. It was first introduced by GeoVG [36], which proposed the first RS visual grounding dataset. MGVLF [37] leverages multiscale visual features and multi-granularity textual embeddings to address the scale variation in RS images. RS referring segmentation is also in its early stages due to the challenges mentioned earlier. It was first introduced by RefSegRS [36], which proposed a new dataset and baseline model. Recently, the transformer-based method RMSIN [16] proposed an adaptive rotated convolution to tackle the issues arising from the scale variation and orientations prevalent in aerial imagery. However, these two types of models have always been studied separately, hindering progress in the field. In this paper, we unify these two tasks within a single framework, allowing them to share data and architecture.
@@ 3.2 Hybrid Supervision
4 refGeo
To address the issue of limited generalization capability in VLMs caused by the relatively small size of existing RS visual grounding datasets, we present a large-scale RS referring expression comprehension dataset, refGeo. It integrates most existing RS visual grounding datasets. The details of each dataset are provided in Tab. 1. Since both GeoChat [9] and VRSBench [13] use DIOR [12] image data, which overlap with DIOR-RSVG [37], we remove the samples corresponding to images that appear in the DIOR-RSVG test and validation sets from the GeoChat and VRSBench training set to prevent data leakage.
Dataset | Format | # Refers | GSD (m) | Image Width | 3D |
[…]
@@
Qwen-VL [1] | 58.76 | 58.65 | 10.59 | 12.99 | 41.75 | 47.38 | 9.53 | 34.24 |
GeoChat [9] | 61.96 | 60.27 | 14.67 | 16.32 | 56.99 | 51.36 | 11.52 | 39.01 |
LLaVA-1.5-7B [15] | 65.98 | 64.46 | 20.95 | 19.98 | 63.76 | 57.17 | 15.05 | 43.91 |
@@
Aerial Vehicle Visual Grounding.
Moreover, we propose a new aerial vehicle visual grounding dataset (AVVG) using unmanned aerial vehicles (UAVs), which presents two key advantages over previous RS visual grounding datasets. 1) 3D Visual Grounding: This dataset leverages UAVs and camera parameters to map pixel coordinates to the camera coordinate system, extending the conventional 2D plane of RS visual grounding tasks into 3D space for the first time. As shown in Fig. 4, part of referring expressions requires the VLM to have spatial reasoning capabilities to accurately locate the object. This extension provides essential data support for further exploration of spatial intelligence in remote sensing, fostering VLMs to perceive 3D space from 2D images projected from the real world, much like humans do. 2) Fine-Grained Attributes: Our dataset is also the first aerial vehicle dataset meticulously annotated with vehicle brands and models. Although visual grounding tasks involving fine-grained vehicle attributes are more challenging, they are more aligned with real-world requirements.
Mask Generation.
To support pixel-level visual grounding tasks, we employ the SAM [8] to generate mask annotations based on existing bounding box annotations. Specifically, we input the bounding box as a prompt along with the image into the SAM-H model, which outputs a mask and its corresponding confidence score. To ensure the quality of the masks, predictions with a confidence score below 0.9 are discarded. In such cases, inspired by SegVG [7], we directly use the bounding box as a mask to provide dense supervision signals. For masks with confidence scores above 0.9, following the idea of RSMIN [16], we apply the morphological dilation algorithm to fix the mask. As a result, we constructed the refGeo dataset, which provides annotations in three forms: HBB, OBB, and segmentation mask.
@@ 5.1 Implementation Details
5.2 Referring Expression Comprehension
@@ Results.
Method | Color | Type | Brand | AVG |
LLaVA-1.5-7B | 1.39 | 1.29 | 0.54 | 1.07 |
GeoGround (N=16) | 9.75 | 5.60 | 5.44 | 6.93 |
5.4 Generalized REC
Settings.
We present an RS Generalized REC benchmark based on AVVG, which differs from standard REC in that one referring expression may correspond to multiple objects. We adhere to the VOC evaluation protocols [6] and evaluate our methods using the AP@0.5 metric.
Results.
Tab. 4 compares the performance of GeoGround and LLaVA-1.5 when using color, type, and brand as referring conditions. The results validate the effectiveness of our hybrid supervision approach and indicate that this task is more challenging than standard REC.
5.5 Referring Expression Segmentation
@@
Tab. 5 demonstrates that GeoGround exhibits superior performance in the pixel-level RS vision grounding task. Unlike the other VLMs, GeoGround does not require the introduction of an additional mask decoder, as it inherently possesses segmentation capabilities. Moreover, we attempt to use SAM [8] to refine the coarse masks generated by GeoGround, which allowed GeoGround to achieve results that match the performance of the best RS referring segmentation model [16]. See Appendix for more details.
Method | Acc@0.5 (Mask) | mIoU | ||
@@
5.6 Visual Understanding
Model | VQA | Image Captioning | REC | |
Acc | METEOR | ROUGE-L | Acc@0.5 | |
LLaVA-1.5-7B | 70.3 | 34.1 | 40.7 | 44.1 |
GeoGround (N=16) | 69.8 | 33.8 | 40.8 | 65.7 |
Settings.
Results.
Tab. 6 compares the visual understanding performance of LLaVA-1.5 and GeoGround, both fine-tuned with the VRSBench [13] versatile instruction set. These results indicate that GeoGround’s performance on VQA and image captioning tasks fluctuates slightly compared to LLaVA-1.5, but shows significant improvement in the REC task. This confirms that our approach enhances region-level understanding without compromising the holistic image comprehension capabilities of VLMs.
@@ 5.7 Ablation Study
Method | DownSample | Resolution | Length | Acc@0.5 |
HBB | - | - | - | 58.74 |
@@
Method | Format | Range | Acc@0.5 | Acc@0.7 |
GeoChat | [, , , , ] | (0, 90] | 26.56 | 8.24 |
polygon | [, ,…, , | - | 26.93 | 11.26 |
oc | [, , , , ] | (0, 90] | 31.29 | 13.61 |
le90 | [, , , , ] | [-90, 90) | 35.78 | 15.43 |
le135 | [, , , , ] | [-45, 135) | 35.39 | 15.56 |
Angle Representation of OBB.
Tab. 11 compares 5 common OBB angle representations in the rotated object detection community [44]. The results indicate that different angle representations affect the visual grounding capability of VLM. Specifically, the angle representation used by GeoChat [9] exhibits poor performance, whereas the two long-edge representations, le90 and le135, demonstrate superior results. Ultimately, we selected le90 as the default angle representation for GeoGround.
@@ 6 Conclusion
Supplementary Material
1 Additional Visual Grounding Signal Details
1.1 Type of Signals used in VLMs.
Tab. 12 summarizes several visual grounding signals currently employed in the VLM domain, all of which use the signal textualization paradigm to convert visual grounding tasks into text generation tasks. To the best of our knowledge, we are the first to convert masks of RS objects into text sequences successfully.
Model | Signal | Density | Resolution | Text Length |
Kosmos-2 [22] | HBB | Sparse | 1/32 | 2 tokens |
Shikra [3] | HBB | Sparse | 1/1000 | 4 numbers |
GeoChat [9] | OBB | Medium | 1/100 | 5 numbers |
EarthGPT [39] | Polygon | Medium | 1/100 | 8 numbers |
Text4Seg [11] | Mask | Dense | 1/N | NN phases |
GeoGround | Mask | Dense | 1/N | NN numbers |
1.2 Format of HBB Signals.
Tab. 13 presents the numerical ranges and representation precision of several common HBB encoding formats in current grounded VLMs. Additionally, we calculated the number of tokens required by the LLaVA tokenizer to encode a single object using these formats. Format #1, adopted by GeoChat [9], requires more tokens than #2 for the same precision. Since the tokenizer requires more tokens to encode decimal values, Format #3, which normalizes coordinates to 1 and retains three decimal places, results in the highest token count. We recommend using #4 or #5, corresponding to the InternVL [5] and Qwen-VL [1] formats, respectively, as they offer greater precision while requiring fewer tokens. Reducing the number of tokens needed to encode a single object not only lowers the training cost for VLMs but also accelerates inference. Ultimately, we selected #4 as the default HBB format for GeoGround.
# | Range | Precision | Format | #Token |
1 | [0,100) | 100 | 1220 | |
2 | [0, 100) | 100 | 1014 | |
3 | [0,1) | 1000 | 1826 | |
4 | [0, 1000) | 1000 | 1018 | |
5 | [0, 1000) | 1000 | 1018 |
1.3 Resolution of Signals.
Fig. 7 compares the disappearance rates of objects across 5 remote sensing visual grounding datasets when different quantization resolutions are applied. It can be observed that the datasets, ranked by the proportion of small targets from highest to lowest at a quantization resolution of 16, are RSVG, AVVG, GeoChat, DIOR-RSVG, and VRSBench. This also explains why the VLM scores are relatively low on the RSVG and AVVG datasets. When the quantization resolution is set to 32, the object disappearance rate remains below 20% for all datasets. Thus, Matrix4Seg achieves satisfactory performance when the resolution is set to 32. At a quantization resolution of 100, GeoChat still has a disappearance rate of 12.65%, indicating the presence of some very small targets in the GeoChat dataset.
2 Additional Data Details
2.1 Mask Generation.
The released refGeo dataset contains manually annotated HBBs or OBBs. However, the masks used in the experiments are generated through an automated method, and thus their quality cannot be fully guaranteed. Although we set a score threshold, errors in the masks still occur. Interestingly, despite the noise in these mask data, they still validate the effectiveness of the hybrid supervision training proposed in this paper. In future work, we plan to manually refine these automatically generated masks in refGeo to provide more accurate referring expression segmentation samples for training.
2.2 Creation of AVVG.
Existing open-source RS datasets [34, 12] often lack sensor metadata and provide limited target attributes. Consequently, these datasets can only support the formulation of simple visual grounding problems, such as locating a object in the upper-left corner of an image. To develop a spatially aware remote sensing dataset, we use unmanned aerial vehicles to collect data from scratch. This approach ensures comprehensive access to sensor parameters, allowing inversion of the three-dimensional coordinates of image targets based on imaging geometry principles. Data collection is divided into two parts: aerial imagery and ground video.
Aerial Imagery.
All aerial images in AVVG are collected with a small UAV platform, DJI Mini3, between 10-16 September 2023, in Shanghai. The dataset consists of 4K high-resolution RS images from 11 distinct scenes, captured at 9 different above-ground levels and 3 pitch angles. This implies that these RS images possess different spatial resolutions and perspectives. In addition, the collected images cover a variety of weather scenarios, such as sunny, cloudy, and rainy days, along with different lighting conditions.
Ground Video.
We record ground videos from the same areas to facilitate accurate annotation of vehicle brands and models. Specifically, we select time slots with relatively low vehicular mobility, avoiding rush hours and meal times. Additionally, to mitigate the vehicle mismatch between drone images and ground videos caused by vehicle entry and exit, we capture two sets of ground videos before and after the drone captures aerial photos. This ensures that vehicles entering or exiting the scene halfway through the capture are recorded in the videos. However, there are instances where vehicles pass through the scene briefly, leading to cases where they are not captured in either video. In such situations, we mask these vehicles with a black mask in the images to ensure that all visible vehicles have fully known attributes. Due to privacy concerns, ground videos will not be released.
Derivation of Coordinate System Transformation.
The transformation between the pixel coordinate system and the image coordinate system can be represented by an affine matrix, as follows:
where represents the pixel size of the sensor. and denote the origin offsets, with the origin of the image coordinate system typically located at the image’s top-left corner. Given the pixel coordinates of a certain point, its corresponding image coordinates can be calculated as follows:
The transformation from the camera coordinate system to the image coordinate system is a conversion from three-dimensional to two-dimensional coordinates. Assuming the focal length of the camera is , then we have
where denotes the depth of the point, which can be obtained by a depth camera (binocular or structured light). Because the drone camera we are using cannot provide depth information, we need to find another way.
When the ground satisfies the ground plane assumption, given the AGL of the drone and the pitch angle of the camera, the ground plane equation in the camera coordinate system is as follows:
The equation of the line connecting the camera origin to the projection point on the pixel plane in the camera coordinate system is given by:
Substituting the line equation into the ground plane equation yields:
Substituting back into the line equation yields:
Fig. 8 visualizes the 3D bounding box of a vehicle. The implementation process is as follows: First, the 3D coordinates of the vehicle’s center point are calculated in the camera coordinate system. Then, based on the orientation information and the vehicle’s length, width, and height, the 3D coordinates of its eight corner points are computed. Finally, these points in the 3D coordinate system are projected back into the 2D-pixel coordinate system. The visualization results further validate the correctness of this coordinate transformation. With the 3D bounding box of each vehicle available, we can formulate visual grounding problems that require the model to have spatial reasoning abilities to solve. For example, which vehicle is the farthest from the camera? Which vehicle has the highest height?
To preserve the spatial mapping between camera coordinates and pixel coordinates, we refrained from cropping the 4K images to increase the dataset size, as is commonly done in most remote sensing datasets. Since the problems involve fine-grained attributes of the vehicles, we used the HBB coordinates of the object as the answers to these spatial grounding tasks, rather than the 3D bounding box, to control the difficulty and prevent the tasks from becoming excessively challenging.
Generalized REC Benchmark.
The generalized referring expression comprehension benchmark built using AVVG includes AP@0.5 scores for three fine-grained vehicle attributes. The specific categories for each attribute are listed as follows: Color (13): golden, brown, orange, silver, gray, blue, white, red, black, pink, green, yellow, purple. Type (10): compact, suv, mid-size, subcompact, mpv, sports, full-size, crossover, sedan, micro-size. Brand (52): skoda, buick, audi, toyota, tesla, volkswagen, mercedes-benz, nio, chevrolet, mazda, hyundai, lexus, honda, bmw, ora, geely, byd, jeep, volvo, renault, roewe, ford, nissan, infiniti, peugeot, cadillac, rising, aito, xpeng, haval, lynk&co, mini, citroen, riich, leap, wey, ideal, changan, harvard, denza, chery, kia, porsche, mg, beijing, smart, baojun, aion, wuling, vgv, hongqi, lexus, lincoln.
3 Additional Implementation Details
We employ the LLaVA framework to train LLaVA-1.5 and GeoChat, while the SWIFT [42] framework is used to train the Qwen-VL model. The main experiments cover 3 types of connectors, including MLP (LLaVA-1.5, GeoChat), Pixel Shuffle + MLP (InternVL2) and Cross-attention (Qwen-VL). In addition, the selected VLMs have different input image resolutions: fixed 336336 for LLaVA-1.5, fixed 448448 for Qwen-VL, fixed 504504 for GeoChat, and dynamic input resolution for InternVL2 and Qwen2-VL.
3.1 Details of Training Hyper-parameters.
Tab. 14 presents the training hyperparameters used to train the GeoGround. We adhere primarily to the same settings as GeoChat [9], and these parameters are also consistently applied in other fine-tuned VLMs in this paper. However, we observed that two novel VLMs that support dynamic resolution images, InternVL2 [5] and Qwen2-VL [28], struggled to achieve satisfactory fine-tuning results with the same set of parameters. Unfortunately, we have not yet identified satisfactory training parameters for them.
Param Name | Value | |
Optimizer | Type | AdamW |
Learning rate | 2e-4 | |
Weight decay | 0.0 | |
(0.9, 0.95) | ||
Gradient norm clip | 1.0 | |
Scheduler | Linearly decay | |
Warmup ratio | 0.03 | |
LoRA | Rank | 64 |
Alpha () | 16 | |
Dropout | 0.05 | |
Module | Linear layers of connector and LLMs | |
Training | Trainable #Params. | About 2% of the LLM (7B 160M) |
Numerical precision | FP16 | |
Global batch size | 128 | |
Total epochs | 5 | |
GPUs | V100(32G) 8 |
3.2 Details of Object Detection Training Samples.
We construct a fine-grained vehicle object detection benchmark leveraging the metadata from the test split of our proposed AVVG dataset. Additionally, to assess performance on remote sensing object detection, we create a dataset consisting of 200k samples by combining data from DOTA [34], DIOR [12], FAIR1M [25], and the train split of AVVG. Specifically, we first cropped the original images from the FAIR1M and DOTA datasets into 512512 patches, while retaining the original image sizes for the DIOR and AVVG datasets. We then used the object detection annotations from these datasets to generate training samples, where queries were formulated based on the categories present in each image. Specifically, if an image contains objects from three categories, we constructed three image-text sample pairs accordingly. These data are used to enhance GeoGround’s basic visual perception capabilities.
3.3 Implementation of Mask Refiner.
Due to resolution limitations, the proposed Matrix4Seg currently struggles to segment particularly small targets. To address this issue, we can simply refine the output masks using SAM to obtain more accurate masks. Specifically, we resize the binary matrix output from the model to match the dimensions of the image, generating a coarse mask. We then sample 3 positive points and 3 negative points from the mask. The sampling region for the positive points is the intersection between the coarse mask and the predicted HBB, while the sampling region for the negative points is the difference between the coarse mask and the predicted HBB. Subsequently, the positive and negative points, along with the predicted HBB, are used as prompts and input into SAM. If no positive or negative sample points are available, only the predicted HBB is used as the prompt for SAM. Since SAM is training-free, it can be regarded as a post-processing method. As shown in Tab. 5, the precision of the masks can be further improved after refinement.
3.4 Visual Instruction Templates
Fig. 10 lists all the instruction templates used during the training of GeoGround. During training, a query is randomly selected from a set of five for each iteration.
4 Additional Quantitative Results
4.1 More Results on REC.
DIOR-RSVG | RSVG | GeoChat | VRSBench | AVVG | Acc@0.5 |
[…]
Tab. 15 shows that as more supervised fine-tuning samples are added to the training set, the Acc@0.5 metric, computed using HBBs on the DIOR-RSVG test set, exhibits a continuous upward trend for the LLaVA-1.5-7B model. This indicates that the scaling law remains effective, suggesting that further increasing the number of visual grounding samples in the RS domain is still a promising avenue.
DIOR | DOTA | FAIR1M | AVVG | Acc@0.5 |
[…]
AVVG | DIOR | DOTA | FAIR1M | Overall AP@0.5 |
[…]
Furthermore, we conducted an in-depth investigation into the impact of object detection on the REC task. As shown in Tab. 16, incorporating object detection samples during training can also improve the REC performance of the model. An interesting observation is that, although DIOR-RSVG is constructed using images from the DIOR dataset for the visual grounding task, incorporating object detection samples from the DIOR dataset only improves the results by approximately 1%. However, when object detection samples from the DOTA dataset are added, whose images come from a different source, the performance increases by more than 10%. This indicates that the improvement in the performance of the REC task through object detection is not only due to data leakage. Since the AVVG data set is designed for fine-grained vehicle attribute detection and only contains vehicles as target objects, it does not contribute to the performance improvement in the DIOR-RSVG data set.
4.2 More Results on Generalized REC.
Tab. 17 shows that adding Generalized REC training samples based on other object detection datasets can also improve the model’s performance on AVVG.
5 Additional Qualitative Results
See more qualitative results of the REC task in Figs. 11 and 12. See more qualitative results of the REC with OBB task in Figs. 13, 14 and 15. See more qualitative results of the RES task in Figs. 16, 17, 18, 19 and 20.

### v3

# GeoGround: A Unified Large Vision-Language Model for Remote Sensing Visual Grounding

@@ Abstract
Remote sensing (RS) visual grounding aims to use natural language expression to locate specific objects (in the form of the bounding box or segmentation mask) in RS images, enhancing human interaction with intelligent RS interpretation systems. Early research in this area was primarily based on horizontal bounding boxes (HBBs), but as more diverse RS datasets have become available, tasks involving oriented bounding boxes (OBBs) and segmentation masks have emerged. In practical applications, different targets require different grounding types: HBB can localize an object’s position, OBB provides its orientation, and mask depicts its shape. However, existing specialized methods are typically tailored to a single type of RS visual grounding task and are hard to generalize across tasks. In contrast, large vision-language models (VLMs) exhibit powerful multi-task learning capabilities but struggle to handle dense prediction tasks like segmentation. This paper proposes GeoGround, a novel framework that unifies support for HBB, OBB, and mask RS visual grounding tasks, allowing flexible output selection. Rather than customizing the architecture of VLM, our work aims to elegantly support pixel-level visual grounding output through the Text-Mask technique. We define prompt-assisted and geometry-guided learning to enhance consistency across different signals. Experimental results show that GeoGround demonstrates strong performance across four RS visual grounding tasks, matching the performance of specialized methods on multiple benchmarks. Code available at https://github.com/zytx121/GeoGround.
@@ 1 Introduction
To support GeoGround training and promote the development of visual RS grounding, we introduce refGeo, a large-scale RS visual grounding instruction-following dataset. refGeo consolidates four existing visual grounding datasets from RS [27, 38, 7, 14] and introduces a new aerial vehicle visual grounding dataset.
@@
We introduce refGeo, the largest RS visual grounding instruction-following dataset, consisting of 161k image-text pairs and 80k RS images.
@@ 2 Related Work
Compared to multimodal tasks like image captioning [41, 15], text-image retrieval [20], and visual question answering (VQA) [18] in RS, research on referring detection is a novel task with limited research. It was first introduced by GeoVG [37], which proposed the first RS visual grounding dataset. MGVLF [38] leverages multiscale visual features and multi-granularity textual embeddings to address the scale variation in RS images. LQVG [10] propose a language query-based Transformer framework for RSVG. LPVA [13] achieves precise attention on referred objects by adjusting visual features with progressive attention. RS referring segmentation is also in its early stages due to the challenges mentioned earlier. It was first introduced by RefSegRS [37], which proposed a new dataset and baseline model. Recently, the transformer-based method RMSIN [17] proposed an adaptive rotated convolution to tackle the issues arising from the scale variation and orientations prevalent in aerial imagery. FIANet [11] proposed a fine-grained image-text alignment module to better discriminative multimodal representation. However, these two types of models have always been studied separately, hindering progress in the field. In this paper, we unify these two tasks within a single framework, allowing them to share data and architecture.
@@ 3.2 Hybrid Supervision
3.3 Dataset
To address the issue of limited generalization capability in VLMs caused by the relatively small size of existing RS visual grounding datasets, we present a large-scale RS referring expression comprehension dataset, refGeo. It integrates most existing RS visual grounding datasets. The details of each dataset are provided in Tab. 1. Since both GeoChat [7] and VRSBench [14] use DIOR [12] image data, which overlap with DIOR-RSVG [38], we remove the samples corresponding to images that appear in the DIOR-RSVG test and validation sets from the GeoChat and VRSBench training set to prevent data leakage. Moreover, we propose a new aerial vehicle visual grounding dataset using unmanned aerial vehicles.
| Dataset | Format | # Refers | GSD (m) | Image Width |
| RSVG [27] | HBB | 5,505 | 0.244.88 | 1,024 |
| DIOR-RSVG [38] | HBB | 27,133 | 0.530 | 800 |
| GeoChat [7] | OBB | 63,883 | 0.30.8 | 6001,024 |
| VRSBench [14] | OBB | 38,689 | 0.130 | 512 |
| Aerial vehicle dataset (Ours) | OBB | 26,465 | 0.0070.04 | 4,000 |
| refGeo | Mixed | 161,675 | 0.00730 | 5124,000 |
@@ 3.3 Dataset
| Qwen-VL (sft) | 58.76 | 58.65 | 10.59 | 12.99 | 41.75 | 47.38 | 9.53 | 34.24 |
| GeoChat (sft) | 61.96 | 60.27 | 14.67 | 16.32 | 56.99 | 51.36 | 11.52 | 39.01 |
| LLaVA-1.5-7B (sft) | 65.98 | 64.46 | 20.95 | 19.98 | 63.76 | 57.17 | 15.05 | 43.91 |
@@ 4 Experiments
4.1 Referring Expression Comprehension (REC)
@@ Results.
4.3 Referring Expression Segmentation (RES)
@@
Tab. 4 demonstrates that GeoGround exhibits superior performance in the pixel-level RS vision grounding task. GeoGround (N=32) achieve better results than directly using HBB to prompt SAM even without relying on SAM. Moreover, we attempt to use HBB and coarse mask to prompt SAM [6], which allowed GeoGround to achieve results that match the performance of the best RS referring segmentation model [17]. See appendix for more details.
| Method | Acc@0.5 | mIoU | ||
@@ 4.4 Ablation Study
| Method | DownSample | Resolution | Length | Acc@0.5 (HBB) |